# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ IV Examples of Topological Phonons: 3D Bulk Materials
In 2018, Zhang et al. (2018a) recognized that the solution to this heavy calculation challenge could be adapted from electronic systems. To simplify electronic surface state calculations, tight-binding models based on localized Wannier functions are typically employed. This approach allows topological surface states to be determined by extending the model Hamiltonian and iteratively solving the Green’s function of a semi-infinite system. Zhang et al. (2018a) observed that the second-order force constant matrix parameters, derived from real-space atomic displacement forces, exhibit formal similarities to $p$-orbital hopping terms in electronic systems. This analogy enables the extension of first-principles tight-binding methods from electronic structure calculations to phonon spectra. Consequently, this methodology can be applied to predict topological phonon materials and compute their surface states with significantly reduced computational cost.
@@
Figure 22: (a) Conventional Weyl phonon. (b-c) Three types of double-Weyl phonons with $|C|=2$, where (b) is the twofold quadratic Weyl point carrying Chern numbers of ±$\pm 2$ for each band, (c) is the threefold spin-1 Weyl point with Chern numbers of 0, ±$\pm 2$ for each band, and (d) is the fourfold charge-2 Dirac point exhibiting Chern numbers of ±$\pm 2$ for each phonon band pair. Zhang et al. (2018a) refer to the quadratic Weyl, spin-1 Weyl, and charge-2 Dirac phonons as double-Weyl phonons. Adapted from Ref. Zhang et al. (2018a).
Figure 23: Crystal structure, phonon spectra, and topological phonons in the $M$Si family ($M$ = Fe, Mn, Co, Re, Ru). (a) Crystal structure of the $M$Si family. (b) Bulk and (001) surface Brillouin zones. (c) Phonon dispersion of FeSi along high-symmetry lines in the rillouin zone. The threefold acoustic phonons are spin-1 double Weyl phonons with Chern numbers of 0 and $\pm$2. Orange and purple boxes highlight the spin-1 Weyl phonon at $\Gamma$ and the charge-2 Dirac phonon at R, respectively. (d) The $\Gamma$ point hosts a spin-1 Weyl phonon. (e) The R point exhibits a charge-2 Dirac phonon. (f) Double-helicoid surface states, described by the Weierstrass elliptic function. The spin-1 Weyl phonon (red cone) corresponds to a double zero point, while the charge-2 Dirac phonon (purple cone) represents a double pole. Two yellow sheets depict topological surface states rotating around these double-Weyl phonons. Adapted from Ref. Zhang et al. (2018a).
@@ IV.1 Topological Phonons with 0D Degeneracies
In addition to the case described by Eq. (34) with $N=2$, double Weyl phonons can emerge in a three-band Hamiltonian $H(\bm{k})=\bm{k}\cdot\bm{L}$, where $\bm{L}$ represents the spin-1 rotation generators. The eigenvalues of $L_{k}=\bm{L}\cdot\bm{k}=\hbar,-\hbar,0$ correspond to Chern numbers $+2,-2,0$, doubling those of the spin-1/2 Weyl phonon. This motivates the term spin-1 Weyl phonon, as illustrated in Fig. 22 (c).
@@ IV.1.1 B-20 Type Materials: Double Weyl Phonons
The first realizations of spin-1 Weyl phonons and charge-2 Dirac phonons were theoretically predicted in the phonon spectra of the $M$Si family ($M$ = Fe, Co, Mn, Re, Ru), as shown in Fig. 23, with experimental confirmation following in the same year. $M$Si family belongs to the B-20 structure, a simple cubic crystal structure with space group $P2_{1}3$ (#198), as shown in Fig. 23 (a). Figure 23 (b) is the BZ and surface BZ (green plane) along (001) direction. Using FeSi as a representative example, Zhang et al. (2018a) identified two types of double Weyl phonons in the $M$Si family. The phonon spectrum shown in Fig. 23 (c) reveals that the three acoustic branches form a spin-1 Weyl phonon. Notably, in 3D chiral crystals, all acoustic phonons exhibit spin-1 Weyl character, where the longitudinal acoustic (LA) phonon carries a zero Chern number, and the two transverse acoustic (TA) phonons have Chern numbers of $\pm$ 2. However, the topological surface states of the spin-1 Weyl phonon remain unobservable due to the absence of a bulk gap between the LA phonon branches.
@@
In the optical phonon spectra of FeSi, all threefold-degenerate modes at the $\Gamma$ point are spin-1 Weyl phonons, while all fourfold-degenerate modes at the R point are charge-2 Dirac phonons, as highlighted by orange and purple boxes in Figs. 23 (c)–(e). The spin-1 Weyl phonons are stabilized by chiral cubic symmetry, whereas the charge-2 Dirac phonon is protected by both crystalline symmetries and time-reversal symmetry $\mathcal{T}$. Zhang et al. (2018a) offered an intuitive perspective to understand the charge-2 Dirac phonon at R. $M$Si family materials possesses twofold screw rotations along each axis, such as $\{C_{2x}|(\frac{1}{2},\frac{1}{2},0)\}$ along $x$ axis, and the other two along $y$ and $z$ axes. These operations anti-commute with each other and satisfy $C_{2x/y/z}^{2}=-1$ at R. Since such behavior is analogous to half-integer spin rotation, all irreducible representations at R must have even dimensions. Moreover, R is a time-reversal-invariant momentum and preserves $\mathcal{T}$, the screw rotations must also commute with $\mathcal{T}$, enforcing real matrix representations. The smallest allowed representation under these constraints is 4D, forcing all bands at R to form charge-2 Dirac phonons, a direct consequence of the interplay between crystalline symmetries and $\mathcal{T}$.
@@
According to the “bulk-surface correspondence”, two surface states must connect these double Weyl phonons, topologically equivalent to helicoid noncompact Riemann sheets Fang et al. (2016). This allows the surface state dispersion near the Weyl phonon projections to be mapped to Riemann surfaces of analytic functions, treating the surface momentum as a complex variable. As an illustrative example, consider the highest optical branch in FeSi ($\sim$14.5 THz). Figures 23 (d) and (e) show zoomed-in dispersions of the spin-1 Weyl phonon at $\Gamma$ (monopole charge +2) and the charge-2 Dirac phonon at R (monopole charge --2), respectively. In the $M$Si family, Weyl phonons carry Chern numbers of $C=\pm 2$, making their (001) surface state dispersion topologically equivalent to the winding phase of the Weierstrass elliptic function (Weierstrass-$\wp$). Specifically, the function has an order-$C$ zero at $\bm{k_{-}}=(0,0)$, and order-$C$ pole at $\bm{k_{+}}=(\pi,\pi)$:
@@
where $z$ represents the planar momentum relative to the Weyl phonon projection, defined as $z\equiv{}k-(1+i)\pi$. As illustrated in Fig. 23 (f), the spin-1 Weyl phonon (red cone) projects to the center of the (001) surface Brillouin zone, while the charge-2 Dirac phonon (blue cone) projects to the Brillouin zone corner. The double-helicoid surface states of the $M$Si family are depicted as yellow sheets. The isoenergetic contour reveals surface arcs (green curves) with distinct topological features. Unlike conventional Weyl materials where surface arcs are open and terminate at bulk Weyl projections, the $M$Si family exhibits a single non-contractible loop that rotates around both double Weyl phonons as the energy contour varies.
@@
It’s worthy to note that all the twofold degeneracies at $\Gamma$ are another type of Weyl phonons with $C=\pm 4$, referred to as twofold quadruple Weyl phonon Zhang et al. (2020b); Li et al. (2021a), as mentioned in Sec. II.2.1. In solids, there are three distinct mechanisms to realize Weyl phonons with a monopole charge of 4, i.e., double spin-1 Weyl phonon, spin-3/2 Weyl phonon and twofold quadruple Weyl phonon, as shown in Figs. 24 (a), (b) and (c), respectively.
@@
The sixfold double spin-1 Weyl node consists of two identical spin-1 Weyl phonons, with its monopole charge arising from the two lowest bands, as shown in Fig. 24 (a). The spin-3/2 Weyl phonon is described by a four-band Hamiltonian of $H(\bm{k})=\bm{k}\cdot\bm{L}$, with $\bm{L}$ representing the spin-3/2 rotation generators. This yields Chern numbers of +1, +3, $-1$, $-3$ for each band. Regarding twofold Weyl phonons, crystals can host three additional types with Chern numbers ±$\pm 1$, ±$\pm 2$ and ±$\pm 3$, respectively. The twofold quadruple Weyl phonon with Chern numbers of ±$\pm 3$ remained unrealized for an extended period (before 2020) due to the oversight of chiral cubic symmetry in topological band theory analyses. We will discuss the twofold Weyl quadruple phonon in detail in Sec. VII.
@@
Xiao et al. pointed out that some kinds of nodal surface can carry integer-valued topological charge similar to Weyl monopoles, i.e. a topological phase transition can occur when the topological charge of the nodal surface changes by simultaneously emitting or absorbing integer numbers of Weyl points Xiao and Fan (2017). In realistic materials, crystals with two-fold screw rotation symmetry $S_{2\alpha}$ ($\alpha=x,y,z$) can host nodal surface phonons at $k_{\alpha}=\pi$ when combined with time-reversal symmetry Liu et al. (2021d); Xie et al. (2021, 2022). Consider a lattice invariant under the nonsymmorphic operation $S_{2x}=\{C_{2x}|[\frac{1}{2}00]\}$. Since $(S_{2x})^{2}=\{E|[100]\}$ represents a primitive lattice translation along $x$ (assuming lattice constant $a$ = 1), Bloch’s theorem gives $(S_{2x})^{2}=e^{ik_{x}}$. At the Brillouin zone boundary $k_{x}=\pi$, the anti-unitary operator $S_{2x}\mathcal{T}$ satisfies $(S_{2x}\mathcal{T})^{2}=(S_{2x})^{2}=-1$, enforcing Kramers-like degeneracy across all bands and consequently generating a nodal surface phonon.
@@
Figure 24: (a–c) Three kinds of Weyl phonons with the monopole charge of 4. (a) Sixfold double spin-1 Weyl phonon, which is a composite of two identical spin-1 Weyl phonons. (b) Fourfold spin-3/2 Weyl phonon with Chern numbers of $+1$, $+3$, $-1$, and $-3$ for each band. (c) Twofold quadruple Weyl phonon with Chern numbers of $+4$ and $-4$ for each band. (d–f) Three types of twofold Weyl phonons with distinct Chern numbers that can exist in crystals, with Chern numbers of ±$\pm 1$, ±$\pm 2$ and ±$\pm 3$, respectively. Adapted from Ref. Zhang et al. (2020b).
@@ IV.1.2 Face-centered Silicon: Triple Point Phonon
$n_{W}=\oint_{C}\frac{\mathrm{d}l}{4\pi i}\leavevmode\nobreak\ \textrm{Tr}\left[\sigma_{z}H_{2}^{-1}\partial_{l}H_{2}\right]=\pm 2.$ (69)
The winding number $n_{W}$ can be defined for any closed loop $C$ encircling the $k_{z}$-axis at fixed $k_{z}$. A nonzero $n_{W}$ directly reflects the topological protection of the nodal structure and is further linked to the Euler characteristic of the band crossing, as demonstrated in Park et al. (2021).
The topology of TPs can also be understood by its topological phase transitions via symmetry breaking terms. When time-reversal symmetry $\mathcal{T}$ is broken while rotational symmetry $C_{n}$ is preserved, an additional term $H^{\prime}_{2}=\delta_{T}\sigma_{z}$ modifies the effective Hamiltonian $H_{2}$. This splits each TP into Weyl point pairs, giving rise to double Fermi-arc-like surface states Liu et al. (2022). On the other hand, if the rotation symmetry $C_{n}$ is broken (e.g., by external strain) while $\mathcal{T}$ remains intact, a perturbation $H^{\prime\prime}_{2}=\delta_{S}\sigma_{x}$ emerges. Unlike the $\mathcal{T}$-breaking case, $H^{\prime\prime}_{2}$ preserves line degeneracy but fragments the nexus into a set of nodal lines, whose positions are determined by
@@
with $A$, $B$, and $C$ being $\bm{k}$-independent constants. Here, $\omega_{0}$ represents the frequency at $\bm{k}=0$, which vanishes for acoustic branches due to the acoustic sum rule, resulting in a linear phonon dispersion near the zone center. This three-band crossing constitutes a topological acoustic triple point (TATP) Park et al. (2021). For optical branches, $\omega_{0}$ remains finite, leading to quadratic $\bm{k}=0$-dependence of the dispersion and the formation of a quadratic contact triple point (QCTP). Remarkably, the QCTP hosts extended topological surface phonon states that persist across the entire Brillouin zone Zhong et al. (2021).
Figure 26: Triple point phonons in diamond-structure silicon. (a) Diamond-structure silicon with space group #227. (b) Phonon band structure of silicon with phononic TPs formed by band crossings between longitudinal optical (LO) and transverse optical (TO) branches along $\Gamma$-X. (c) Distribution of topological nodal lines (red solid) and TPs in momentum space. (d) Schematics of TPs and double-fermi-arc-like surface states under time-reversal symmetry (TRS). By breaking the TRS, the TPs are splitted into pairs of WPs which support single-fermi-arc surface states. (e) Under [001] uniaxial strain, TPs transform into interlocking nodal Hopf links. Adapted from Ref. Liu et al. (2022).
To date, numerous materials hosting phononic TPs and QCTPs have been theoretically predicted through first-principles calculations. Li et al. (2018) identified three-component Weyl phonons in WC-type compounds (TiS, ZrSe, HfTe) within the THz frequency range. Subsequent work by Singh et al. (2018) further explored these systems, emphasizing their thermoelectric properties. In CsCl, Park et al. (2021) demonstrated acoustic TPs with distinct topological invariants: the longitudinal mode carries a Skyrmion number, while the transverse modes are characterized by an Euler number. Phononic QCTPs have been reported in cubic materials such as Zr3Ni3Sb4 Zhong et al. (2021), Ta3Sn Yang et al. (2022), etc.

### v2

@@ IV Examples of Topological Phonons: 3D Bulk Materials
In 2018, zhang2018double recognized that the solution to this heavy calculation challenge could be adapted from electronic systems. To simplify electronic surface state calculations, tight-binding models based on localized Wannier functions are typically employed. This approach allows topological surface states to be determined by extending the model Hamiltonian and iteratively solving the Green’s function of a semi-infinite system. zhang2018double observed that the second-order force constant matrix parameters, derived from real-space atomic displacement forces, exhibit formal similarities to $p$-orbital hopping terms in electronic systems. This analogy enables the extension of first-principles tight-binding methods from electronic structure calculations to phonon spectra. Consequently, this methodology can be applied to predict topological phonon materials and compute their surface states with significantly reduced computational cost.
In realistic solids, additional care should be paid when diagnosing the topological index of phonon band structure. Sometimes the bare force-constant approach fails. One example is the LO-TO splitting effect in polar semiconductors. In polar crystals, the long-range macroscopic electric field induced by the LO mode can induce a nonlocal term in the dynamical matrix as $\mathsf{D}(\bm{q})=\mathsf{D}^{\textrm{local}}(\bm{q})+\mathsf{D}^{\textrm{nonlocal}}(\bm{q})$ where the local term $\mathsf{D}^{\textrm{local}}(\bm{q})$ is ordinary dynamical matrix based on the second-order force constants given by Eq. (4) and the nonlocal term $\mathsf{D}^{\textrm{nonlocal}}(\bm{q})$ is given by Gonze1997Apr ; Wang2010Apr
$\mathsf{D}^{\textrm{nonlocal}}_{i\alpha,j\beta}(\bm{q})=\left\{\begin{array}[]{cl}\displaystyle\frac{4\pi}{V_{\textrm{uc}}\sqrt{M_{i}M_{j}}}\frac{(\mathsf{Z}_{i}\cdot\bm{q})_{\alpha}(\mathsf{Z}_{j}\cdot\bm{q})_{\beta}}{\bm{q}\cdot\mathsf{\epsilon}_{\infty}\cdot\bm{q}}&(\bm{q}\rightarrow 0)\\ 0&(\bm{q}=0)\end{array}.\right$. (72)
$V_{\textrm{uc}}$ is the volume of unit cell, and $M_{i,j}$ and $\mathsf{Z}_{i,j}$ refer to the atomic mass and Born effective charge matrix of $i,j$-th atom, respectively. $\epsilon_{\infty}$ refers to the dielectric matrix in the high-frequency limit. The nonlocal term split the degeneracy of LO and TO modes at $\bm{q}=0$ and is non-analytical. xu2024catalog show that this non-analytical correction (NAC) can induce additional band inversion which may change the topological index of phonon spectrum. Dy2O3 is such an example. Without NAC, the lowest 11 phonon bands of Dy2O3 form a topologically nontrivial subset which is totally gapped from the other bands. However, if NAC is taken into account the band order at $\Gamma$ is further switched which makes the subset topologically trivial.
Another interesting aspect of polaritonic effect is about the topological phonon polaritons Karzig2015Jul ; Guddala2021Oct ; Hu2020Jun ; Wang2023Mar . When coupling a transverse optical (TO) phonon branch with a photon mode, the “flat” TO mode can hybridize with the photon branch to open a band gap as shown in Fig. 19. In this hybrid phonon-photon system, the “phonon” and “photon” sectors can be viewed as two different orbitals of the effective Hamiltonian and the band inversion is clearly seen in Fig. 19(a) between $k=0$ and $k\rightarrow\infty$. The effective Hamiltonain can be expressed as
$H=\left(\begin{array}[]{cc}\omega_{\textrm{phonon}}(\bm{k})&\Delta\\ \Delta^{\dagger}&\omega_{\textrm{photon}}(\bm{k})\end{array}\right),$ (73)
[…]
Figure 19: Schematic band structure of phonon polariton. (a) If we treat the “phonon” part and “photon” part as two different orbitals of the hybrid Hamiltonian, the band structure shows a clear band inversion between $k=0$ and $k\rightarrow\infty$. (b) Topological polaritonic edge states, adapted from Karzig2015Jul .
@@
Figure 20: (a) Conventional Weyl phonon. (b-c) Three types of double-Weyl phonons with $|C|=2$, where (b) is the twofold quadratic Weyl point carrying Chern numbers of $\pm 2$ for each band, (c) is the threefold spin-1 Weyl point with Chern numbers of 0, $\pm 2$ for each band, and (d) is the fourfold charge-2 Dirac point exhibiting Chern numbers of $\pm 2$ for each phonon band pair. zhang2018double refer to the quadratic Weyl, spin-1 Weyl, and charge-2 Dirac phonons as double-Weyl phonons. Adapted from Ref. zhang2018double .
Figure 21: Crystal structure, phonon spectra, and topological phonons in the MMSi family ($M$ = Fe, Mn, Co, Re, Ru). (a) Crystal structure of the MMSi family. (b) Bulk and (001) surface Brillouin zones. (c) Phonon dispersion of FeSi along high-symmetry lines in the rillouin zone. The threefold acoustic phonons are spin-1 double Weyl phonons with Chern numbers of 0 and $\pm$2. Orange and purple boxes highlight the spin-1 Weyl phonon at $\Gamma$ and the charge-2 Dirac phonon at R, respectively. (d) The $\Gamma$ point hosts a spin-1 Weyl phonon. (e) The R point exhibits a charge-2 Dirac phonon. (f) Double-helicoid surface states, described by the Weierstrass elliptic function. The spin-1 Weyl phonon (red cone) corresponds to a double zero point, while the charge-2 Dirac phonon (purple cone) represents a double pole. Two yellow sheets depict topological surface states rotating around these double-Weyl phonons. Adapted from Ref. zhang2018double .
@@ IV.1 Topological Phonons with 0D Degeneracies
In addition to the case described by Eq. (39) with $N=2$, double Weyl phonons can emerge in a three-band Hamiltonian $H(\bm{k})=\bm{k}\cdot\bm{L}$, where $\bm{L}$ represents the spin-1 rotation generators. The eigenvalues of $L_{k}=\bm{L}\cdot\bm{k}=\hbar,-\hbar,0$ correspond to Chern numbers +2,2,0+2,-2,0, doubling those of the spin-1/2 Weyl phonon. This motivates the term spin-1 Weyl phonon, as illustrated in Fig. 20 (c).
@@ IV.1.1 B-20 Type Materials: Double Weyl Phonons
The first realizations of spin-1 Weyl phonons and charge-2 Dirac phonons were theoretically predicted in the phonon spectra of the MMSi family ($M$ = Fe, Co, Mn, Re, Ru), as shown in Fig. 21, with experimental confirmation following in the same year. MMSi family belongs to the B-20 structure, a simple cubic crystal structure with space group $P2_{1}3$ (#198), as shown in Fig. 21 (a). Figure 21 (b) is the BZ and surface BZ (green plane) along (001) direction. Using FeSi as a representative example, zhang2018double identified two types of double Weyl phonons in the MMSi family. The phonon spectrum shown in Fig. 21 (c) reveals that the three acoustic branches form a spin-1 Weyl phonon. The acoustic phonons inherently exhibit threefold degeneracy at the $\Gamma$ point. This degeneracy is lifted in chiral crystals. Thus, in 3D chiral crystals, all acoustic phonons exhibit spin-1 Weyl character, where the longitudinal acoustic (LA) phonon carries a zero Chern number, and the two transverse acoustic (TA) phonons have Chern numbers of $\pm$ 2. However, the topological surface states of the spin-1 Weyl phonon remain unobservable due to the absence of a bulk gap between the LA phonon branches.
@@
In the optical phonon spectra of FeSi, all threefold-degenerate modes at the $\Gamma$ point are spin-1 Weyl phonons, while all fourfold-degenerate modes at the R point are charge-2 Dirac phonons, as highlighted by orange and purple boxes in Figs. 21 (c)–(e). The spin-1 Weyl phonons are stabilized by chiral cubic symmetry, whereas the charge-2 Dirac phonon is protected by both crystalline symmetries and time-reversal symmetry $\mathcal{T}$. zhang2018double offered an intuitive perspective to understand the charge-2 Dirac phonon at R. MMSi family materials possesses twofold screw rotations along each axis, such as $\{C_{2x}|(\frac{1}{2},\frac{1}{2},0)\}$ along $x$ axis, and the other two along $y$ and $z$ axes. These operations anti-commute with each other and satisfy $C_{2x/y/z}^{2}=-1$ at R. Since such behavior is analogous to half-integer spin rotation, all irreducible representations at R must have even dimensions. Moreover, R is a time-reversal-invariant momentum and preserves $\mathcal{T}$, the screw rotations must also commute with $\mathcal{T}$, enforcing real matrix representations. The smallest allowed representation under these constraints is 4D, forcing all bands at R to form charge-2 Dirac phonons, a direct consequence of the interplay between crystalline symmetries and $\mathcal{T}$.
@@
According to the “bulk-surface correspondence”, two surface states must connect these double Weyl phonons, topologically equivalent to helicoid noncompact Riemann sheets fang2016topological . This allows the surface state dispersion near the Weyl phonon projections to be mapped to Riemann surfaces of analytic functions, treating the surface momentum as a complex variable. As an illustrative example, consider the highest optical branch in FeSi ($\sim$14.5 THz). Figures 21 (d) and (e) show zoomed-in dispersions of the spin-1 Weyl phonon at $\Gamma$ (monopole charge +2) and the charge-2 Dirac phonon at R (monopole charge -2), respectively. In the MMSi family, Weyl phonons carry Chern numbers of $C=\pm 2$, making their (001) surface state dispersion topologically equivalent to the winding phase of the Weierstrass elliptic function ($\wp$). Specifically, the function has an order-$C$ zero at $\bm{k_{-}}=(0,0)$, and order-$C$ pole at $\bm{k_{+}}=(\pi,\pi)$:
@@
where $z$ represents the planar momentum relative to the Weyl phonon projection, defined as $z\equiv{}k-(1+i)\pi$. As illustrated in Fig. 21 (f), the spin-1 Weyl phonon (red cone) projects to the center of the (001) surface Brillouin zone, while the charge-2 Dirac phonon (blue cone) projects to the Brillouin zone corner. The double-helicoid surface states of the MMSi family are depicted as yellow sheets. The isoenergetic contour reveals surface arcs (green curves) with distinct topological features. Unlike conventional Weyl materials where surface arcs are open and terminate at bulk Weyl projections, the MMSi family exhibits a single non-contractible loop that rotates around both double Weyl phonons as the energy contour varies.
@@
It’s worthy to note that all the twofold degeneracies at $\Gamma$ are another type of Weyl phonons with $C=\pm 4$, referred to as twofold quadruple Weyl phonon zhang2020twofold ; li2021observation , as mentioned in Sec. II.3.1. In solids, there are three distinct mechanisms to realize Weyl phonons with a monopole charge of 4, i.e., double spin-1 Weyl phonon, spin-3/2 Weyl phonon and twofold quadruple Weyl phonon, as shown in Figs. 22 (a), (b) and (c), respectively.
@@
The sixfold double spin-1 Weyl node consists of two identical spin-1 Weyl phonons, with its monopole charge arising from the two lowest bands, as shown in Fig. 22 (a). The spin-3/2 Weyl phonon is described by a four-band Hamiltonian of $H(\bm{k})=\bm{k}\cdot\bm{L}$, with $\bm{L}$ representing the spin-3/2 rotation generators. This yields Chern numbers of +1, +3, 1-1, 3-3 for each band. Regarding twofold Weyl phonons, crystals can host three additional types with Chern numbers $\pm 1$, $\pm 2$ and $\pm 3$, respectively. The twofold quadruple Weyl phonon with Chern numbers of $\pm 3$ remained unrealized for an extended period (before 2020) due to the oversight of chiral cubic symmetry in topological band theory analyses. We will discuss the twofold Weyl quadruple phonon in detail in Sec. VII.
@@
Xiao et al. pointed out that some kinds of nodal surface can carry integer-valued topological charge similar to Weyl monopoles, i.e. a topological phase transition can occur when the topological charge of the nodal surface changes by simultaneously emitting or absorbing integer numbers of Weyl points xiao2017topologically . In realistic materials, crystals with two-fold screw rotation symmetry $S_{2\alpha}$ ($\alpha=x,y,z$) can host nodal surface phonons at $k_{\alpha}=\pi$ when combined with time-reversal symmetry Liu2021Jul_NS ; Xie2021Oct ; Xie2022Feb . Consider a lattice invariant under the nonsymmorphic operation $S_{2x}=\{C_{2x}|[\frac{1}{2}00]\}$. Since $(S_{2x})^{2}=\{E|[100]\}$ represents a primitive lattice translation along $x$ (assuming lattice constant $a$ = 1), Bloch’s theorem gives $(S_{2x})^{2}=e^{ik_{x}}$. At the Brillouin zone boundary $k_{x}=\pi$, the anti-unitary operator $S_{2x}\mathcal{T}$ satisfies $(S_{2x}\mathcal{T})^{2}=(S_{2x})^{2}=-1$, enforcing Kramers-like degeneracy across all bands and consequently generating a nodal surface phonon.
@@
Figure 22: (a–c) Three kinds of Weyl phonons with the monopole charge of 4. (a) Sixfold double spin-1 Weyl phonon, which is a composite of two identical spin-1 Weyl phonons. (b) Fourfold spin-3/2 Weyl phonon with Chern numbers of +1+1, +3+3, 1-1, and 3-3 for each band. (c) Twofold quadruple Weyl phonon with Chern numbers of +4+4 and 4-4 for each band. (d–f) Three types of twofold Weyl phonons with distinct Chern numbers that can exist in crystals, with Chern numbers of $\pm 1$, $\pm 2$ and $\pm 3$, respectively. Adapted from Ref. zhang2020twofold .
@@ IV.1.2 Face-centered Silicon: Triple Point Phonon
$n_{W}=\oint_{C}\frac{\mathrm{d}l}{4\pi i}\penalty 10000\ \textrm{Tr}\left[\sigma_{z}H_{2}^{-1}\partial_{l}H_{2}\right]=\pm 2.$ (77)
The winding number $n_{W}$ can be defined for any closed loop $C$ encircling the $k_{z}$-axis at fixed $k_{z}$. A nonzero $n_{W}$ directly reflects the topological protection of the nodal structure and is further linked to the Euler characteristic of the band crossing, as demonstrated in park2021topological ; peng2022phonons ; peng2022multigap .
The topology of TPs can also be understood by its topological phase transitions via symmetry breaking terms. When time-reversal symmetry $\mathcal{T}$ is broken while rotational symmetry $C_{n}$ is preserved, an additional term $H^{\prime}_{2}=\delta_{T}\sigma_{z}$ modifies the effective Hamiltonian $H_{2}$. This splits each TP into Weyl point pairs, giving rise to double Fermi-arc-like surface states liu2022ubiquitous . On the other hand, if the rotation symmetry $C_{n}$ is broken (e.g., by external strain) while $\mathcal{T}$ remains intact, a perturbation $H^{\prime\prime}_{2}=\delta_{S}\sigma_{x}$ emerges. Unlike the $\mathcal{T}$-breaking case, $H^{\prime\prime}_{2}$ preserves line degeneracy but fragments the nexus into a set of nodal lines, whose positions are determined by
@@
with $A$, $B$, and $C$ being $\bm{k}$-independent constants. Here, $\omega_{0}$ represents the frequency at $\bm{k}=0$, which vanishes for acoustic branches due to the acoustic sum rule, resulting in a linear phonon dispersion near the zone center. This three-band crossing constitutes a topological acoustic triple point (TATP) park2021topological . For optical branches, $\omega_{0}$ remains finite, leading to quadratic $\bm{k}=0$-dependence of the dispersion and the formation of a quadratic contact triple point (QCTP). Remarkably, the QCTP hosts extended topological surface phonon states that persist across the entire Brillouin zone Zhong2021Aug .
Figure 24: Triple point phonons in diamond-structure silicon. (a) Diamond-structure silicon with space group #227. (b) Phonon band structure of silicon with phononic TPs formed by band crossings between longitudinal optical (LO) and transverse optical (TO) branches along $\Gamma$-X. (c) Distribution of topological nodal lines (red solid) and TPs in momentum space. (d) Schematics of TPs and double-fermi-arc-like surface states under time-reversal symmetry (TRS). By breaking the TRS, the TPs are splitted into pairs of WPs which support single-fermi-arc surface states. (e) Under [001] uniaxial strain, TPs transform into interlocking nodal Hopf links. Adapted from Ref. liu2022ubiquitous .
To date, numerous materials hosting phononic TPs and QCTPs have been theoretically predicted through first-principles calculations. li2018coexistent identified three-component Weyl phonons in WC-type compounds (TiS, ZrSe, HfTe) within the THz frequency range. Subsequent work by singh2018topological further explored these systems, emphasizing their thermoelectric properties. In CsCl, park2021topological demonstrated acoustic TPs with distinct topological invariants: the longitudinal mode carries a Skyrmion number, while the transverse modes are characterized by an Euler number. Phononic QCTPs have been reported in cubic materials such as Zr3Ni3Sb4 Zhong2021Aug , Ta3Sn Yang2022Mar , etc.